# Reliability, and Global Privacy

### v3

@@ IV Toward Fair and Secure Aggregation: Fair-SecCoGC
It can be verified that $\forall k:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},6\bm{I}_{D})$ and that $\sum_{k=1}^{5}\bm{N}_{k}^{t}=\mathbf{0}$, ensuring fair privacy preservation while achieving exact recovery of the global model.
V Local Mutual Information Privacy (LMIP)
This section examines LMIP of each local model update in SecCoGC at all levels in the learning process. At the $t$-th round, both clients and the server are aware of the previous global model $\bm{\Theta}^{t}$, knowing $\Delta\bm{\Theta}_{k}^{t,I}$ is equivalent to knowing $\bm{\Theta}_{k}^{t,I}$. For simplicity, we omit the index $I$ in the notation below.
V-1 Peer-to-Peer LMIP
This section aims to quantify the maximum information leakage during the gradient-sharing phase, when the best secret keys are used. The clients are assumed to be honest but curious, they are honest because they faithfully perform their role in the system. But they can be interested in learning sensitive information about other clients. We assume that each secret key $\bm{N}_{k}^{t}\in\mathbb{R}^{D}$ satisfies a total power constraint, such that $\operatorname{Tr}(\operatorname{Cov}[\bm{N}_{k}^{t}])\leq D\lambda^{2}$. Similarly, we assume that each local update $\Delta\bm{\Theta}_{k}^{t}\in\mathbb{R}^{D}$ satisfies a total power constraint of $D\zeta^{2}$, such that $\operatorname{Tr}(\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])\leq D\zeta^{2}$. The local model updates $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ are assumed to be mutually independent since local training is performed independently by each client.
Theorem V.1 (Peer-to-Peer LMIP).
In both SecCoGC and Fair-SecCoGC, the transmission from client $k$ to $m$ satisfy $\mu_{1}$-LMIP, where
$\mu_{1}=(1-p_{m,k})\frac{D}{2}\log\left(1+\frac{\zeta^{2}}{\lambda^{2}}\right).$ (27)
Proof.
The maximal information leakage provided by the best privacy noise can be formulated as follows,
$\adjustlimits{\sup}_{\Delta\bm{\Theta}_{k}^{t}}{\inf}_{\bm{N}_{k}^{t}}\mathrm{I}(\Delta\bm{\Theta}_{k}^{t};\tau_{m,k}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})|\tau_{m,k})$ (28a)
$=$ $\adjustlimits{\sup}_{\Delta\bm{\Theta}_{k}^{t}}{\inf}_{\bm{N}_{k}^{t}}(1-p_{m,k})\mathrm{I}(\Delta\bm{\Theta}_{k}^{t};\tau_{m,k}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})|\tau_{m,k}=1)$ (28b)
$=$ $\adjustlimits{\sup}_{\Delta\bm{\Theta}_{k}^{t}}{\inf}_{\bm{N}_{k}^{t}}(1-p_{m,k})\left(\mathrm{h}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})-\mathrm{h}(\bm{N}_{k}^{t})\right)$ (28c)
$=$ $\sup_{\Delta\bm{\Theta}_{k}^{t}}(1-p_{m,k})\left(\mathrm{h}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})-\mathrm{h}\left(\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})\right)\right)$ (28d)
$=$ $(1-p_{m,k})\left(\mathrm{h}\left(\mathcal{N}(\mathbf{0},(\lambda^{2}+\zeta^{2})\bm{I}_{D})\right)-\mathrm{h}\left(\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})\right)\right)$ (28e)
$=$ $(1-p_{m,k})\frac{D}{2}\log\left(1+\frac{\zeta^{2}}{\lambda^{2}}\right),$ (28f)
where (28b) owes to fact that if $\tau_{m,k}=0$, then client $m$ learns nothing about client $k$, i.e., $\mathrm{I}(\Delta\bm{\Theta}_{k}^{t};\tau_{m,k}(\Delta\bm{\Theta}_{k}^{t}+\bm{N}_{k}^{t})|\tau_{m,k}=0)=0$, (28c) is due to the independence between $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ and $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$, (28d) and (28e) follows from the fact that, under a total power (trace) constraint, the maximum differential entropy is achieved by a Gaussian vector with a diagonal covariance matrix whose diagonal entries are equal. ∎
The above results lead to two key conclusions. First, Fair-SecCoGC not only ensure fair privacy protection among clients, but also produces optimal keys $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$ under a total power constraint of $D\lambda^{2}$, in the sense that it offers the smallest information leakage. Therefore, in the analysis below, we employ these optimal secret keys for all $k$. Second, any other input source than a Gaussian vector with a diagonal covariance matrix $\zeta^{2}\bm{I}_{D}$ results in lower information leakage.
V-2 LMIP at Relay
This section examines the information leakage in the partial sums when a client acts as a central relay. The secret keys $\left\{\bm{N}_{k}^{t}\right\}_{k=1}^{K}$ are correlated across $k$, due to the coordinated nature of secret keys in secure aggregation. The server is assumed to follow the protocol faithfully but may attempt to infer sensitive information about $\Delta\bm{\Theta}_{k}^{t}$. In the strictest case, where the server has access to the full connectivity realizations $\bm{\mathcal{T}}:=[\tau_{m,k}^{t}]_{K\times K}$ among clients, the information leakage for each local update, contained in the observed partial sums (complete or incomplete), is characterized below.
Theorem V.2 (LMIP in Masked Partial Sums).
In Fair-SecCoGC, the partial sums computed by client $k$, in protecting client $k_{1}$’s contribution, satisfy $\mu_{2}$-LMIP, where
$\mu_{2}=(1-p_{k,k_{1}})\frac{D}{2}\cdot$
$\left(\log\left((2\pi e)g_{k,k_{1}}^{2}\zeta^{2}\right)+\log\left(\frac{\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}}{\lVert\bm{\Lambda}_{k}^{t,-k_{1}}\bm{A}\rVert^{2}}\right)\right).$ (29)
Proof.
The server receives the partial sum $\bm{S}_{k}^{t}$ from client $k$ and wishes to infer sensitive information about client $k_{1}$. The information leakage can be expressed by
$\mathrm{I}\left(\Delta\bm{\Theta}_{k_{1}}^{t};\bm{S}_{k}^{t}|\bm{\mathcal{T}},\bm{G}\right)$
$=$ $\mathrm{I}\left(\Delta\bm{\Theta}_{k_{1}}^{t};\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$=$ $\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$\hskip 22.76219pt-\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\Delta\bm{\Theta}_{k_{1}}^{t},\bm{\mathcal{T}},\bm{G}\right)$
$=$ $\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$\hskip 22.76219pt-\mathrm{h}\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\left(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t}\right)\bigg{|}\bm{\mathcal{T}},\bm{G}\right),$ (30)
where $\bm{\Lambda}_{k}^{t}(m)=\tau_{k,m}g_{k,m}$ is the actually received coefficient by client $k$ from client $m\in[K]$. The (30) follows that $\bm{\Lambda}_{k}^{t}(k_{1})$ is fully determined by $\bm{\mathcal{T}}$, and $\bm{G}$ and that $\Delta\bm{\Theta}_{k_{1}}^{t}$ is independent of $\{\Delta\bm{\Theta}_{k}^{t}\}_{k\neq k_{1}}$, $\bm{\mathcal{T}}$, and $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$. Due to Guassian distribution and independence between $\{\Delta\bm{\Theta}_{k}^{t}\}_{k=1}^{K}$ and $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$, the differential entropy in (30) becomes separable,
$\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)\hskip-2.84526pt+\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)$
- $\mathrm{h}\left(\hskip-2.84526pt\sum_{m\neq k_{1}}\hskip-2.84526pt\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)\hskip-2.84526pt-\mathrm{h}\hskip-2.84526pt\left(\hskip-2.84526pt\sum_{m\neq k_{1}}\hskip-2.84526pt\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right).$ (31)
Due to Guassian distribution and independence among $\Delta\bm{\Theta}_{k}^{t}$ for all $k$ and independent communications, the sum can be taken out, (31) further becomes
$\sum_{m=1}^{K}\mathrm{h}\left(\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)+\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$-\hskip-5.69054pt\sum_{m\neq k_{1}}\hskip-2.84526pt\mathrm{h}\left(\bm{\Lambda}_{k}^{t}(m)\Delta\bm{\Theta}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\hskip-2.84526pt\right)\hskip-2.84526pt-\mathrm{h}\hskip-2.84526pt\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$=\mathrm{h}\left(\bm{\Lambda}_{k}^{t}(k_{1})\Delta\bm{\Theta}_{k_{1}}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)+\mathrm{h}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right)$
$-\mathrm{h}\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\bigg{|}\bm{\mathcal{T}},\bm{G}\right).$ (32)
Let $\bm{\Lambda}_{k}^{t}=[\bm{\Lambda}_{k}^{t}(1),\cdots,\bm{\Lambda}_{k}^{t}(K)]$ stores all the actually received coefficients at client $k$, it further holds that
$\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}=\bm{\Lambda}_{k}^{t}\begin{bmatrix}\bm{N}_{1}^{t}\\ \vdots\\ \bm{N}_{K}^{t}\end{bmatrix}=\bm{\Lambda}_{k}^{t}\bm{A}\cdot\begin{bmatrix}\bm{Z}_{1}^{t}\\ \vdots\\ \bm{Z}_{L}^{t}\end{bmatrix}$. (33)
Thus, the variance of the aggregated correlated Gaussian privacy noise $\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{k}^{t}$ is given by
$\nu_{k}^{t}\triangleq\mathrm{var}\left(\sum_{m=1}^{K}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\right)=\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}\bm{I}_{D}.$ (34)
Let $\bm{\Lambda}_{k}^{t,-k_{1}}$ denote $\bm{\Lambda}_{k}^{t}$ but with $\bm{\Lambda}_{k}^{t}(k_{1})=0$. Similarly, we get
$\nu_{k}^{t,-k_{1}}\triangleq\mathrm{var}\left(\sum_{m\neq k_{1}}\bm{\Lambda}_{k}^{t}(m)\bm{N}_{m}^{t}\right)=\lVert\bm{\Lambda}_{k}^{t,-k_{1}}\bm{A}\rVert^{2}\bm{I}_{D}.$ (35)
Finally, the right side of(32) becomes
$\frac{D}{2}(1-p_{k,k_{1}})\log\left((2\pi e)g_{k,k_{1}}^{2}\zeta^{2}\right)\frac{D}{2}\log\left(\frac{\lVert\bm{\Lambda}_{k}^{t}\bm{A}\rVert^{2}}{\lVert\bm{\Lambda}_{k}^{t,-k_{1}}\bm{A}\rVert^{2}}\right).$ (36)
[…]
V-3 LMIP at the Server
To better showcase the conclusion, we assume $\operatorname{Cov}[\Delta\bm{\Theta}_{k}^{t}])=\mathrm{Diag}(\bm{I}_{D}\zeta^{2})$.
Theorem V.3 (LMIP in Server Aggregation).
In secure aggregation, the server can reconstruct the exact global model $\sum_{m=1}^{K}w_{m}\Delta\bm{\Theta}_{m}^{t}$ without learning about each local update $\Delta\bm{\Theta}_{k}^{t}$. The global model satisfies $\mu_{3}$-LMIP, where
$\mu_{3}=\frac{D}{2}\log\left(1+\frac{w_{k}^{2}}{\sum_{m\neq k}w_{m}^{2}}\right).$ (37)
In the special case, where $w_{k}=\frac{1}{M}$, $\mu_{3}=\frac{D}{2}\log\left(1+\frac{1}{M-1}\right)$.
Proof.
The information leakage about $\Delta\bm{\Theta}_{k}^{t}$ in the exact global model is given by
$\mathrm{I}\left(\Delta\bm{\Theta}_{k}^{t};\sum_{m=1}^{K}w_{m}\Delta\bm{\Theta}_{m}^{t}\bigg{|}\{w_{m}\}_{m=1}^{K}\right)$
$=$ $\mathrm{h}\hskip-2.84526pt\left(\sum_{m=1}^{K}\hskip-2.84526ptw_{m}\Delta\bm{\Theta}_{m}^{t}\bigg{|}\{w_{m}\}_{m=1}^{K}\hskip-2.84526pt\right)\hskip-2.84526pt-\hskip-1.42262pt\mathrm{h}\hskip-2.84526pt\left(\hskip-2.84526pt\sum_{m\neq k}\hskip-2.84526ptw_{m}\Delta\bm{\Theta}_{m}^{t}\bigg{|}\{w_{m}\}_{m=1}^{K}\hskip-2.84526pt\hskip-1.42262pt\right)$
$=$ $\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},\sum_{m=1}^{K}w_{m}^{2}\zeta^{2}\bm{I}_{D}\right)\right)-\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},\sum_{m\neq k}w_{m}^{2}\zeta^{2}\bm{I}_{D}\right)\hskip-2.84526pt\right)$
$=$ $\frac{D}{2}\log\left(1+\frac{w_{k}^{2}}{\sum_{m\neq k}w_{m}^{2}}\right).$ (38)
Remark V.1.
Lemma V.3 provides a key insight into privacy behavior in secure aggregation. As the weight $w_{k}$ increases, the information leakage $\mu_{3}$ also increases. This indicates that, for model parameters of comparable scale, greater importance leads to weaker privacy guarantees the global model. Interestingly, when $w_{k}=\frac{1}{K}$, i.e., all clients are equally important, we observe that $\lim_{K\rightarrow\infty}\mu_{3}\rightarrow 0$. However, this vanishing leakage does not imply asymptotically perfect privacy in large-scale systems. On the contrary, it stems from the fact that $\lim_{K\rightarrow\infty}\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},K\cdot\frac{1}{K^{2}}\zeta^{2}\bm{I}_{D}\right)\right)=\lim_{K\rightarrow\infty}\mathrm{h}\left(\mathcal{N}\left(\mathbf{0},(K-1)\cdot\frac{1}{K^{2}}\zeta^{2}\bm{I}_{D}\right)\right)=$-$\infty$, which implies no privacy.
VI Local Differential Privacy (LDP)
This section examines the model privacy by LDP across all levels in Fair-SecCoGC. We begin by presenting the relevant definitions, adapted to our context for clarity and readability.
Definition VI.1 ($(\epsilon,\delta)$-LDP).
Definition VI.2 (Guassian Mechanism[3]).
VI-1 Peer-to-Peer LDP
The local model updates are masked locally prior to transmission as in (6), i.e., $f(\Delta\bm{\Theta}_{k}^{t})=\Delta\bm{\Theta}_{k}^{t}$. Even if client $m$ can decode client $k$’s transmission, it should not confidently learn about $\Delta\bm{\Theta}_{k}^{t}$.
Theorem VI.1 (Peer-to-Peer LDP).
Given the secret key $\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$, $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, and $\forall\delta_{m,k}^{(1)}\in(0,1]$, the masked transmission from client $k$ to client $m$ via unreliable link $\mathrm{Ber}(1-p_{m,k})$ is $(\epsilon_{m,k}^{(1)},(1-p_{m,k})\delta_{m,k}^{(1)})$-differentially private, where
$\epsilon_{m,k}^{(1)}=\frac{2R}{\lambda}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (42)
@@ VI-1 Peer-to-Peer LDP
Proofs are provided in [32, Appendix A].  [^2] The supplementary file can be found at https://arxiv.org/abs/2507.07565.
VI-2 LDP at Relay

### v4

@@ IV-B Fair Zero-Sum Noise Construction
In this example, $\forall k:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},6\bm{I}_{D})$ and $\sum_{k=1}^{5}\bm{N}_{k}^{t}=\mathbf{0}$.
In the remainder of this work, the fair zero-sum noises are employed, obtained from Construction IV.3.
Lemma IV.4 (Correlation Among Privacy Noises).
From Construction IV.3, $\forall k\in[K]:\bm{N}_{k}^{t}\sim\mathcal{N}(\mathbf{0},\lambda^{2}\bm{I}_{D})$. In the $d$-th dimension, where $d\in\{1,\cdots,D\}$, the covariance matrix $\mathrm{Cov}(\bm{N}_{1}^{t}(d),\cdots,\bm{N}_{K}^{t}(d))=\bm{A}\bm{A}^{\top}\triangleq\bm{R}$. Consequently, for a pair of secret keys, $\forall k,m\in[K]$,
$\mathrm{Cov}\left(\bm{N}_{k}^{t}(d),\bm{N}_{m}^{t}(d)\right)=\begin{pmatrix}\lambda^{2}&\bm{R}_{k,m}\\ \bm{R}_{m,k}&\lambda^{2}\end{pmatrix}$, (25)
where $\bm{R}_{k,m}$ is the $k,m$-th element of $\bm{R}$. The conditional distribution of $\bm{N}_{k}^{t}$ given $\bm{N}_{m}^{t}=\bm{n}_{m}$ is Gaussian: $\bm{N}_{k}^{t}\mid\bm{N}_{m}^{t}=\bm{n}_{m}\sim\mathcal{N}\!\left(\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$.
V Local Differential Privacy (LDP)
This section examines the model privacy by LDP across all levels in SecCoGC. The relevant definitions, adapted to our context, are presented in Appendix A.
V-A Peer-to-Peer LDP
The local model update is masked by (7) before transmission. Even if client $m$ can decode client $k$’s transmission, it should not confidently learn about $\Delta\bm{\Theta}_{k}^{t}$. However, unlike [16, 4], client $m$ has access to $\bm{N}_{m}^{t}=\bm{n}_{m}$ as auxiliary information when inferring the $k$-th local model, and this auxiliary information is correlated with $\bm{N}_{k}^{t}$. According to Lemma IV.4, this equivalent to adding independent privacy noise $\mathcal{N}\left(\frac{\bm{n}_{m}}{\lambda^{2}}\bm{R}_{m,k},\;(\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m})\cdot\bm{I}_{D}\right)$. This leads us to the following Theorem.
Theorem V.1 (Peer-to-Peer LDP).
Given the secret key $\{\bm{N}_{k}^{t}\}_{k=1}^{K}$ and their covariance matrix $\bm{R}\triangleq\bm{A}\bm{A}^{\top}$ in each dimension, $\Delta\bm{\Theta}_{k}^{t},\Delta\bm{\Theta}_{k}^{t^{\prime}}\in\mathbb{B}_{D}(R)$, and $\forall\delta_{m,k}^{(1)}\in(0,1]$, the masked transmission from client $k$ to client $m$ via unreliable link $\mathrm{Ber}(1-p_{m,k})$ is $(\epsilon_{m,k}^{(1)},(1-p_{m,k})\delta_{m,k}^{(1)})$-differentially private, where
$\epsilon_{m,k}^{(1)}=\frac{2R}{\sqrt{\lambda^{2}-\frac{1}{\lambda^{2}}\bm{R}_{m,k}\bm{R}_{k,m}}}\left[2\log\left(\frac{1.25}{\delta_{m,k}^{(1)}}\right)\right]^{\frac{1}{2}}.$ (26)
@@ V-A Peer-to-Peer LDP
Proofs are provided in Appendix A. ∎
V-B LDP at the Relaying Client
After communication among clients, the relaying client $k$ computes the partial sum $\bm{S}_{k}^{t}$ in (9), based on the received masked local updates in $\mathcal{U}_{k}^{t}$. Let $\bm{S}_{k}^{t,-j}\triangleq\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}(\Delta\bm{\Theta}_{m}^{t}+\bm{N}_{m}^{t})$ denote the partial sum excluding client $j$’s contribution. Let $\bm{M}_{k}^{t}=\sum_{m\in\mathcal{U}_{k}^{t}}g_{k,m}\bm{N}_{m}^{t}$ be the aggregated privacy noises and $\nu_{k}^{t}\bm{I}_{D}=\mathrm{var}(\bm{M}_{k}^{t})$ be its variance. Similarly, define $\bm{M}_{k}^{t,-j}=\sum_{m\in\mathcal{U}_{k}^{t}\setminus\{j\}}g_{k,m}\bm{N}_{m}^{t}$, and its variance $\nu_{k}^{t,-j}\bm{I}_{D}=\mathrm{var}(\bm{M}_{k}^{t,-j})$. Define axillary matrix $\bm{\Lambda}_{k}^{t}=[\bm{\Lambda}_{k}^{t}(1),\cdots,\bm{\Lambda}_{k}^{t}(K)]$ with $\bm{\Lambda}_{k}^{t}(m)=\tau_{k,m}g_{k,m}$, and $\bm{\Lambda}_{k}^{t,-k_{1}}$ denote $\bm{\Lambda}_{k}^{t}$ with its $k_{1}$-th entry set to zero.